# Bayesian variable selection in sample selection models using spike-and-slab priors

### v2

@@ Acknowledgements
Adam Iqbal was supported by the Heilbronn Institute for Mathematical Research via funding from the ESPRC grant “Additional Funding Programme for Mathematical Sciences” (EP/V521917/1).
Appendix A Induced prior for $\rho$
@@
Appendix B Gibbs sampler for sample selection models
@@
Appendix C Further details on the Gibbs samplers
C.1 Class I prior
The prior variance matrix for the outcome equation, $\bm{B}$, corresponds to the $(p+1)\times(p+1)$ diagonal matrix where $B_{1,1}=\eta^{O}v^{O}_{0}$ and $B_{j,j}=\left((1-\gamma^{O}_{j-1})\tau_{0,\beta}^{2}v+\gamma^{O}_{j-1}\tau_{1,\beta}^{2}\right)v^{O}_{j-1}$ (with the mixing variable $v^{O}_{j-1}$ appearing to allow for scale mixture of normal priors). $\bm{A}$ is defined similarly. Define $\bm{W}_{0},\bm{W}_{1},\bm{s^{*}}_{0},\bm{s^{*}}_{1},\bm{y}_{1},\bm{g},\bm{G}$ and $\bm{Z}$ as in Section B. Then we have that
@@ C.2 Class II prior
$d^{*}$ $=d+\frac{\tilde{\rho}^{2}}{2\tau}+\frac{1}{2}\sum_{\{i:s_{i}=1\}}\left(y_{i}-\bm{x}_{i}{{}^{\top}}\bm{\beta}-\tilde{\rho}(s^{*}_{i}-\bm{w}_{i}^{\top}\bm{\alpha})^{2}\right)+\frac{\beta_{0}^{2}}{2\eta^{O}v^{O}_{0}}+\sum_{j=1}^{p}\frac{\beta_{j}^{2}}{2v^{O}_{j}((1-\gamma^{O}_{j})\tau_{0,\beta}^{2}+\gamma^{O}_{j}\tau_{1,\beta}^{2})}\,.$
Appendix D RAND data table
@@
Appendix E Computational times
To test computational times, we recorded the running time of the methods on one replicate from the simulation study. That is, we simulated $\bm{x}_{i}=(x_{i,1},\ldots,x_{i,p})^{\top}$ as i.i.d. normal for $i=1,\ldots,n$, let $\bm{w}=\bm{x}$, then generated bivariate normal errors to obtain $(y_{i},s_{i})^{\top}$ from (1), with $\sigma=1$. We chose $\rho=0.5$ and to avoid convergence issues, $n=1000$. For the spike-and-slab normal, the chain was run for $10,000$ iterations, discarding the first $1,250$ as burn-in, as in the simulation study. We varied $p$ from $25$ to $200$ (and let $q=p$) to see how the running times scale to very high dimensions. The computations were performed on a computer with 16GB of RAM and an AMD Ryzen 5 3600 GPU.
Method $p=25$ $p=50$ $p=100$ $p=200$
Spike-and-slab normal 15.5 39.3 130.7 609.0
Adaptive LASSO 22.1 79.5 454.4 5156.0
Stepwise selection 23.3 82.2 234.4 1469.6
Table 7: Computational times for different methods, for $n=1000$. All times are in seconds.
Appendix F Additional tables for different correlations
@@ F.1 $n=500$
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
F.2 $n=1000$
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix G Class II prior comparison
@@ Appendix H Tables for $n=2000$
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix I Additional tables for sensitivity analysis
I.1 Varying spike variance for ambulatory data
Spike variance scaling ×$\times 1$$\times 0.5$$\times 2$
@@ I.1 Varying spike variance for ambulatory data
Table 20: Results from ambulatory data from varying the slab variances. The first column refers to what $\tau_{0,\alpha}$ and $\tau_{0,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.
I.2 Varying slab for RAND data
Slab variance scaling ×$\times 1$$\times 0.5$$\times 2$
@@ I.2 Varying slab for RAND data
I.3 Varying spike for RAND data
Spike variance scaling ×$\times 1$$\times 0.5$$\times 2$
@@ I.3 Varying spike for RAND data
Table 22: Results from RAND data from varying the slab variances. The first column refers to what $\tau_{0,\alpha}$ and $\tau_{0,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.

### v3

@@ Acknowledgements
Adam Iqbal was supported by the Heilbronn Institute for Mathematical Research via funding from the EPSRC grant “Additional Funding Programme for Mathematical Sciences” (EP/V521917/1).
Appendix A Computational times
To test computational times for the spike-and-slab Gibbs sampler, we recorded the running time of the method on one replicate from the simulation study. That is, we simulated $\bm{x}_{i}=(x_{i,1},\ldots,x_{i,p})^{\top}$ as i.i.d. normal for $i=1,\ldots,n$, let $\bm{w}=\bm{x}$, then generated bivariate normal errors to obtain $(y_{i},s_{i})^{\top}$ from (1), with $\sigma=1$. We chose $\rho=0.5$ and to avoid convergence issues, $n=1000$. For the spike-and-slab normal, the chain was run for $10,000$ iterations, discarding the first $1,250$ as burn-in, as in the simulation study. We varied $p$ from $25$ to $200$ (and let $q=p$) to see how the running times scale to very high dimensions. The computations were performed on a computer with 16GB of RAM and an AMD Ryzen 5 3600 GPU. The reason this was run separately from the rest of the simulation study is because the original simulations were run in parallel and only considered up to $p=50$.
Method $p=25$ $p=50$ $p=100$ $p=200$
Spike-and-slab normal 15.5 39.3 130.7 609.0
Table 4: Computational times for the spike-and-slab prior, for $n=1000$. All times are in seconds.
Appendix B Induced prior for $\rho$
@@
Appendix C Gibbs sampler for sample selection models
@@
Appendix D Further details on the Gibbs samplers
D.1 Class I prior
The prior variance matrix for the outcome equation, $\bm{B}$, corresponds to the $(p+1)\times(p+1)$ diagonal matrix where $B_{1,1}=\eta^{O}v^{O}_{0}$ and $B_{j,j}=\left((1-\gamma^{O}_{j-1})\tau_{0,\beta}^{2}v+\gamma^{O}_{j-1}\tau_{1,\beta}^{2}\right)v^{O}_{j-1}$ (with the mixing variable $v^{O}_{j-1}$ appearing to allow for scale mixture of normal priors). $\bm{A}$ is defined similarly. Define $\bm{W}_{0},\bm{W}_{1},\bm{s^{*}}_{0},\bm{s^{*}}_{1},\bm{y}_{1},\bm{g},\bm{G}$ and $\bm{Z}$ as in Section C. Then we have that
@@ D.2 Class II prior
$d^{*}$ $=d+\frac{\tilde{\rho}^{2}}{2\tau}+\frac{1}{2}\sum_{\{i:s_{i}=1\}}\left(y_{i}-\bm{x}_{i}{{}^{\top}}\bm{\beta}-\tilde{\rho}(s^{*}_{i}-\bm{w}_{i}^{\top}\bm{\alpha})^{2}\right)+\frac{\beta_{0}^{2}}{2\eta^{O}v^{O}_{0}}$
$+\sum_{j=1}^{p}\frac{\beta_{j}^{2}}{2v^{O}_{j}((1-\gamma^{O}_{j})\tau_{0,\beta}^{2}+\gamma^{O}_{j}\tau_{1,\beta}^{2})}\,.$
Appendix E RAND data table
@@
Appendix F Main simulation tables
@@ F.1 $n=500$
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.783 3.176 0.986 0.969 0.808 3.066 0.976 0.980
Laplace 0.645 3.426 0.990 0.935 0.726 3.240 0.983 0.959
Stepwise 0.355 4.002 0.999 0.856 0.363 3.986 0.996 0.857
ALASSO 0.558 3.458 0.982 0.927 0.762 2.919 0.945 0.988
$p=25$
Normal 0.764 3.043 0.964 0.993 0.794 2.964 0.957 0.996
Laplace 0.735 3.128 0.970 0.990 0.775 3.035 0.964 0.994
Stepwise 0.046 6.269 0.997 0.851 0.036 6.159 0.999 0.856
ALASSO 0.420 3.900 0.980 0.956 0.702 2.874 0.927 0.996
$p=50$
Normal 0.741 2.923 0.941 0.998 0.703 2.833 0.920 0.998
Laplace 0.720 3.003 0.947 0.997 0.698 2.914 0.931 0.997
Stepwise 0.001 10.295 0.998 0.845 0.002 10.194 0.997 0.847
ALASSO 0.271 4.570 0.973 0.965 0.537 2.717 0.871 0.998
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
F.2 $n=1000$
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.876 3.138 1.000 0.980 0.929 3.079 1.000 0.989
Laplace 0.799 3.243 1.000 0.965 0.863 3.157 1.000 0.978
Stepwise 0.350 4.013 1.000 0.855 0.360 4.004 1.000 0.857
ALASSO 0.877 3.123 0.999 0.982 0.974 2.980 0.992 1.000
$p=25$
Normal 0.908 3.094 0.999 0.996 0.928 3.066 0.999 0.997
Laplace 0.871 3.145 0.999 0.993 0.896 3.105 0.999 0.995
Stepwise 0.048 6.077 1.000 0.860 0.044 6.018 1.000 0.863
ALASSO 0.650 3.493 1.000 0.978 0.938 3.033 0.995 0.998
$p=50$
Normal 0.919 3.077 0.999 0.998 0.943 3.041 0.997 0.999
Laplace 0.876 3.127 0.999 0.997 0.918 3.079 0.998 0.998
Stepwise 0.002 9.855 1.000 0.854 0.001 9.625 1.000 0.859
ALASSO 0.476 3.833 0.999 0.982 0.938 3.009 0.991 0.999
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix G Additional tables for different correlations
G.1 $n=500$
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
G.2 $n=1000$
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix H Class II prior comparison
@@ Appendix I Tables for $n=2000$
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix J Additional tables for sensitivity analysis
J.1 Varying slab variance for ambulatory data
Slab variance scaling ×$\times 1$$\times 0.25$$\times 4$
PIP Est. S.D. PIP Est. S.D. PIP Est. S.D.
Selection equation
(Intercept) - 1.276 0.038 - 1.269 0.037 - 1.276 0.038
educ 1.000 0.176 0.032 1.000 0.172 0.032 0.996 0.180 0.033
age 0.949 0.108 0.039 0.977 0.108 0.034 0.914 0.108 0.044
income 0.349 0.029 0.042 0.471 0.037 0.044 0.206 0.019 0.037
female 1.000 0.320 0.031 1.000 0.319 0.030 1.000 0.320 0.031
totchr 1.000 0.603 0.055 1.000 0.586 0.052 1.000 0.604 0.055
blhisp 1.000 -0.171 0.029 1.000 -0.169 0.029 1.000 -0.174 0.029
ins 0.571 0.049 0.046 0.693 0.057 0.044 0.388 0.035 0.044
Outcome equation
(Intercept) - 6.564 0.060 - 6.562 0.057 - 6.566 0.059
educ 0.116 0.008 0.018 0.196 0.011 0.021 0.062 0.005 0.014
age 1.000 0.230 0.026 1.000 0.229 0.026 1.000 0.231 0.026
female 1.000 0.158 0.032 1.000 0.158 0.030 1.000 0.157 0.031
totchr 1.000 0.399 0.031 1.000 0.398 0.030 1.000 0.398 0.031
blhisp 0.895 -0.089 0.039 0.952 -0.092 0.033 0.844 -0.085 0.042
ins 0.033 -0.002 0.008 0.079 -0.002 0.010 0.017 -0.001 0.008
$\sigma$ 1.288 0.024 - 1.287 0.023 - 1.289 0.024
$\rho$0.265 0.154 - -0.262 0.145 - -0.271 0.150
J.2 Varying spike variance for ambulatory data
Spike variances scaling ×$\times 1$$\times 0.25$$\times 4$
@@
Table 22: Results from ambulatory data from varying the spike variances. The first column refers to what $\tau_{0,\alpha}$ and $\tau_{0,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.
J.3 Varying slab for RAND data
Slab variance scaling ×$\times 1$$\times 0.25$$\times 4$
@@
J.4 Varying spike for RAND data
Spike variance scaling ×$\times 1$$\times 0.25$$\times 4$
@@
Table 24: Results from RAND data from varying the spike variances. The first column refers to what $\tau_{0,\alpha}$ and $\tau_{0,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.